OpenStudy (anonymous):

@fortune_paul @iPwnBunnies @ganeshie8

OpenStudy (ipwnbunnies):

Yesssss?

OpenStudy (ipwnbunnies):

midpoint fomula = \[(\frac{ \left( x1+x2 \right) }{ 2 }, \frac{ y1+y2 }{ 2 })\]